A 1.50 m long closed glass tube, a so-called Customer tube, has a speaker at one end. The tube is filled with a gas and in addition there are small cork crumbs that show nodes and bends in the standing longitudinal sound waves that arise in the tube. When the sound from the speaker had the frequency 870 Hz, a standing wave with a total of 7 pcs. nodes, including those at the endpoints of the tube. Determine the speed of sound in the gas.
1
Expert's answer
2020-04-17T10:40:14-0400

As per the question,

Total length of the pipe = 1.5m

Frequency of the sound wave = 870 Hz,

Total number of wave = 7

let the wave length of the wave ="(\lambda)"

"7\lambda = 1.5"

"\lambda=\dfrac{1.5}{7}"

Hence, velocity of the wave ="f\lambda=870\times \dfrac{1.5}{7}=180.42 m/sec"